ramen:

Help Plz: How many atoms of just Oxygen in 50.0g of Al(OH)3?

ramen:

This is a dimensional analysis problem but I am stucc on it

24kgold:

Chemical change, paper is changed to CO2 and H2O

ramen:

That has nothing to do with the problem but thanks for trying

24kgold:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen That has nothing to do with the problem but thanks for trying \(\color{#0cbb34}{\text{End of Quote}}\) ywwwwwwww

24kgold:

sorry if it didn't help

ramen:

It's all good

darkknight:

@vocaloid

Vocaloid:

Divide 50.0g by the molar mass of Al(OH)3 to get moles of Al(OH)3 Then multiply that result by 3 since each molecule has 3 oxygen atoms Now you have moles of oxygen atoms. Multiply that by Avogadro’s number to get the # of individual atoms

ramen:

@Vocaloid So I have to find the molar mass of the whole entire AI(OH)3 to start?

Vocaloid:

Yeah

ramen:

Ok thanks for your help

ramen:

Sorry @Vocaloid I'm Just wondering, do you know how I would set this up as a dimensional analysis format?

Vocaloid:

\[\frac{ g~Al(OH)_{3} }{ }\frac{ 1~mole~Al(OH)3 }{ molar~mass~g~Al(OH)_{3} }\frac{ 3~mole~oxygen }{ 1~moleAl(OH)_{3} }\frac{ 6.022*10^{23}~O~atoms }{ 1~mole~oxygen }\]

ramen:

Where would the 50.0 Grams be included? Is that in the very first part?

Vocaloid:

Yes